OpenStudy (anonymous):

Simplify: (its in fraction form by the way) 6xy^-1/2x^-2y^-1 Simplify: ( its in fraction form by the way) (3x^-2y^)-1/(2xy^-2)^0 Simplify: (its in fraction form by the way) (ax^2)^-1/a^-2x^-2

OpenStudy (john_es):

You should use the laws of powers, \[1/x^{-n}=x^n\]\[x^nx^m=x^{n+m}\]